Adenine had a molecular mass of 135.1 and consist of 44.45%C, 3.13%H and 52.82%N by mass. Determine its molecular formular
1
Expert's answer
2020-04-18T07:08:28-0400
  • The molecular mass of carbon in adenine can be calculated by multiplying the molecular mass of adenine by the percentage of carbon:

M(C) = 135.1(g/mol)•(44.45%/100%) = 60.05 g/mol "\to" then we can calculate the number of the carbon in adenine:

"N" (of carbon's atoms) = "M" (summary mass of all carbon's atoms)"\/M" (mass of one carbon's atom) = 60.05(g/mol)/12.01(g/mol) = 5

  • Then, similarly, we first calculate the molecular masses of hydrogen and nitrogen, and then the number of their atoms in adenine:

M(H) = 135.1(g/mol)•(3.13%/100%) = 4.5 g/mol "\to" "N" (of hydrogen's atoms) = "M" (summary mass of all hydrogen's atoms)/"M"(mass of one hydrogen's atom) = 4.5(g/mol)/1(g/mol) = 5 (in rounding)

M(N) = 135.1(g/mol)•(52.82%/100%) = 71.36(g/mol) "\to" "N" (of nitrogen's atoms) = "M" (summary mass of all nitrogen's atoms)/"M"(mass of one nitrogen's atom) = 71.36(g/mol)/14(g/mol) = 5

Answer: the molecular formula of adenine is C5H5N5